OpenStudy (anonymous):

[2.02] Solve y over negative 2+ 2 = 8 −20 20 12 −12

OpenStudy (anonymous):

\[\frac{ y }{2 }+2=8?\]

OpenStudy (wolf1728):

Is it (y / -2) + 2 = 8 ?

OpenStudy (anonymous):

oh i forgot negative, O.O

OpenStudy (anonymous):

y/(-2)+2=8 y/(-2)=6 y=-12